4. Beginning with PV = nRT, derive ρ = PM/(RT). Note M = molecular weight and m = mass.
5. Derive p2=p1exp(-gMdeltaZ/Rt) for an isothermal, perfect gas.
6. List the equations for pressure and temperature of an isentropic perfect gas.
Answer #1

Question 1

PV=nRT

(n=w/M)

P= wRT/VM (w/V=ρ )

P= ρRT/M

ρ= PM/RT

Question 2

From hydrostatic equation

dp+ ρgdZ = 0 (putting value of ρ= PM/RT)

\frac{dp}{p} + \frac{gM dZ}{RT} = 0

\frac{P2}{P1} = exp(\frac{-gM(z2- z1)}{RT})